Session: Feb 2006 (session 1)
Wavelength: 5 cm


To give feedback on this experiment, click

Experiment Feedback

Medicina (orfei at Wed Feb 22 18:00:54 2006) Success :

Torun (K. Borkowski at Thu Feb 23 07:09:09 2006) Failure : ANT
See the feedback for EB031A. The telescope remains frozen in altitude.

Hartebeesthoek (Jonathan Quick at Thu Feb 23 07:44:06 2006) Success :
No known problems

Onsala (Michael Lindqvist at Thu Feb 23 08:39:03 2006) Success :
No known problems.

Westerbork (Tony Foley at Thu Feb 23 10:24:41 2006) Success :
Single dish was used. Please change the coordinates in the vexfile accordingly. No known problems. The LO switch was handled automatically by a script (called from the SNAP via a procedure).

Effelsberg (A. Polatidis at Tue Feb 28 09:36:34 2006) Success :

Jodrell_Bank (Alastair Gunn at Fri Mar 3 17:08:03 2006) Failure : ANT
The MkII telescope was used (Jb2). Offsets were incorrectly set hence sources probably out of telescope beam.

Cambridge (Alastair Gunn at Fri Mar 3 17:08:28 2006) Success :
No known problems.

Noto (P. Cassaro at Tue Mar 7 15:40:49 2006) Success :

EVN experiment feedback facility: entire session

Entire Session

Session: Feb 2006 (session 1)
Wavelengths: 6 3.6 5 18 cm


To give feedback on this entire session, click

Entire Session Feedback

Torun (Kaz Borkowski at Mon Feb 20 11:25:29 2006) Minor failures : PCAL
Tr phase cal system failed to produce signals during all the 6 cm experiments for yet unknown reason.

Westerbork (Tony Foley at Fri Feb 24 15:48:24 2006) Minor failures : ANT POIN
The feedback from the antenna control for SINGLE-DISH observations is unreliable for 20seconds after a source change. We will try to fix this. (Affected only the 5cm experiments -Z.P.)